OpenStudy (anonymous):

I'll FAN AND MEDAL!!! The profit earned by selling a quantity of x + 2 of a product is given by the expression x4 − x3 − 24. What is the average profit per unit of product sold? A) x^3 − 3x^2 + 18 B) x^3 − 3x^2 + 6x − 12 C) x^3 − 3x^2 + 6x − 12/x+2 D) x^3 − 3x^2 − 18 E) x^3 − 3x^2 − 18/x+2

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

b

OpenStudy (freckles):

so we have the profit=x^4-x^3-24 and the selling quantity=x+2 to find the average product per unit sold we do (x^4-x^3-24)/(x+2)

OpenStudy (freckles):

in other words we want to divide x^4-x^3-24 by x+2 you can do this by long division or synthetic division

OpenStudy (anonymous):

oh ok i'll use synthetic division

OpenStudy (freckles):

to find the average profit per unit sold we do (x^4-x^3-24)/(x+2) * (correcting a word I put earlier)

OpenStudy (anonymous):



OpenStudy (anonymous):

start off with this

OpenStudy (freckles):

awesome you put the zero placeholders in

OpenStudy (anonymous):

and @thisotherliz are you sure

OpenStudy (anonymous):